OpenStudy (anonymous):

how do you write an equation in slope intercept form for the line that satisfies each set of conditions. Find a line that passes through (-4,2) parallel to the line whose equation is y=1/2x+5

OpenStudy (anonymous):

y=1/2x+4

OpenStudy (anonymous):

how did you do that?

OpenStudy (anonymous):

It is parallel to y = 1/2x+5. Parallel lines have the same slope. So our new equation is y = 1/2x + b ---I say b because we don't know b yet Now, we just have to put in the -4 x and 2 for y, because this is one point y = 1/2x + b (2) = 1/2(-4) + b 2= -4/2 + b 2 = -2 + b 4 = b. Now we put in 4 for b, and we have our equation of a line! y = 1/2x + b y = 1/2x + 4 Wala!

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

No problem. Glad to help.